# Two-Stage Nuisance Function Estimation for Causal Mediation Analysis

### v1

@@ 5 Analyses
We start by making the observation that the true nuisance functions $\pi_{1}$ and $\pi_{2}$ satisfy conditions in (1) and (2) with equality.
@@
    If $\hat{\pi}_{1}=\pi_{1}$, then
    $\mathbb{E}\left[\left\{\frac{1-A}{\hat{\pi}_{1}(X)}-1\right\}h_{1}(X)\right]=0,~{}~{}~{}~{}~{}~{}\forall h_{1}\in\mathcal{H}_{1}.$
@@
    If $\hat{\pi}_{1}=\pi_{1}$ and $\hat{\pi}_{2}=\pi_{2}$, then
    $\mathbb{E}\left[\left\{\frac{A\hat{\pi}_{2}(M,X)}{1-\hat{\pi}_{1}(X)}-\frac{1-A}{\hat{\pi}_{1}(X)}\right\}h_{2}(M,X)\right]=0,~{}~{}~{}~{}~{}~{}\forall h_{2}\in\mathcal{H}_{2}.$
Suppose conditions in (1) and (2) hold with equality and $\mathcal{H}_{1}=L^{2}(X)$ and $\mathcal{H}_{2}=L^{2}(M,X)$. Then, it is easy to see that $\hat{\pi}_{1}=\pi_{1}$ and $\hat{\pi}_{2}=\pi_{2}$ with probability one. However, when $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are not the corresponding $L^{2}$ spaces, we can have conditions in (1) and (2) hold with equality even though $\hat{\pi}_{1}\neq\pi_{1}$ and $\hat{\pi}_{2}\neq\pi_{2}$. Nevertheless, satisfying conditions in (1) and (2) can be sufficient for controlling the bias of the final estimator of the parameter of interest. We formalize this argument by looking at the structure of the bias of $\hat{\psi}^{mr}$.
@@
The bias of the estimator $\hat{\psi}^{mr}$ satisfies the following equality.
$\mathbb{E}[\hat{\psi}^{mr}]-\psi_{0}$ $=\mathbb{E}\left[\left\{\frac{(1-A)}{\hat{\pi}_{1}(X)}-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\left[\left\{\frac{A\hat{\pi}_{2}(M,X)}{1-\hat{\pi}_{1}(X)}-\frac{(1-A)}{\hat{\pi}_{1}(X)}\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right].$
From the structure of the bias of the estimator $\hat{\psi}^{mr}$, it is evident that if $\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}$ can be well-approximated by the functions in $\mathcal{H}_{1}$, and $\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}$ can be well-approximated by the functions in $\mathcal{H}_{2}$, then we can eliminate the bias of the estimator $\hat{\psi}^{mr}$ even if we have not had a perfect fit of $\pi_{1}$ and $\pi_{2}$. This is in fact the main motivation of our two-stage approach. It substitutes fitting $p(A=1\mid X)$ and $p(M\mid A,X)$, which can be a difficult task, by a much weaker requirement based on balancing.
Importantly, our approach retains the multiple-robustness property, as stated in the following result.
@@
$\hat{\psi}^{mr}$ is unbiased if at least one of the following pairs of nuisance function estimators is correctly specified.
  • $\{\hat{\pi}_{1},\hat{\pi}_{2}\}$
  • $\{\hat{\mu}_{1},\hat{\mu}_{2}\}$
  • $\{\hat{\pi}_{1},\hat{\mu}_{1}\}$
Proposition 4 immediately follows from Proposition 3 and Theorem 1. Figure 2 summarizes the result in Proposition 4.
@@ 5.1 Asymptotic Normality
    $\big{\|}\frac{p(A=0\mid X)}{\hat{\pi}_{1}(X)}-1\big{\|}_{2}\|\mu_{2}(X)-\hat{\mu}_{2}(X)\|_{2}=o_{p}(n^{-1/2})$, and
    $\big{\|}\frac{p(A=1\mid M,X)\hat{\pi}_{2}(M,X)}{1-\hat{\pi}_{1}(X)}-\frac{p(A=0\mid M,X)}{\hat{\pi}_{1}(X)}\big{\|}_{2}\|\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\|_{2}=o_{p}(n^{-1/2})$.
Define $O:=(X,A,M,Y)$, and for any estimates of nuisance functions $\{\tilde{\pi}_{1},\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2}\}$, let
$m(O;\tilde{\pi}_{1}$ $,\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2}):=\tilde{\mu}_{2}(X)+\frac{(1-A)}{\tilde{\pi}_{1}(X)}\{\tilde{\mu}_{1}(M,X)-\tilde{\mu}_{2}(X)\}+\frac{A\tilde{\pi}_{2}(M,X)}{1-\tilde{\pi}_{1}(X)}\{Y-\tilde{\mu}_{1}(M,X)\}.$
We have the following result regarding the asymptotic normality of the estimator $\hat{\psi}^{mr}$.
@@
Suppose nuisance functions and the parameter of interest are estimated on disjoint samples with sample sizes of the same order as the total sample size. Under Assumption 2, the estimator $\hat{\psi}^{mr}$ is asymptotically normal and satisfies
$\sqrt{n}(\hat{\psi}^{mr}-\psi_{0})=\sqrt{n}\mathbb{E}_{n}\left[m(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]+o_{p}(1).$
As a corollary of Theorem 2, one can use the influence function of the proposed estimator to obtain confidence intervals for the parameter of interest.
@@ 6 Conclusion
In estimating a functional of interest that involves nuisance functions, if one naively aims to perfectly fit the nuisance functions, the resulting estimator of the parameter of interest can be sensitive to even small errors in the nuisance function estimators. Therefore, in order to estimate a functional, it is important to investigate the role of nuisance functions involved in that functional in the structure of the bias. Based on this vision, we proposed a new approach for estimating the nuisance functions of the mediation functional when the influence function-based estimator is used for the parameter of interest. Our proposed estimators were based on the approach of balancing that can target the bias of the parameter of interest more directly and also can lead to an estimator which is less sensitive to error. We demonstrated that this approach retains the multiple-robustness property. Moreover, we provided conditions on our nuisance function estimators sufficient for asymptotic normality of the influence function-based estimator of the mediation functional.
@@ Appendix A Proofs
$\arg\max_{h_{1}}\mathbb{E}\left[\left\{\frac{1-A}{\pi_{1}(X)}-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]$
$=\arg\min_{h_{1}}\mathbb{E}\left[\left\{\frac{1-A}{\pi_{1}(X)}-1\right\}^{2}+\frac{1}{4}h^{2}_{1}(X)-\left\{\frac{1-A}{\pi_{1}(X)}-1\right\}h_{1}(X)\right]$
$=\arg\min_{h_{1}}\mathbb{E}\left[\left(\frac{1}{2}h_{1}(X)-\left\{\frac{1-A}{\pi_{1}(X)}-1\right\}\right)^{2}\right]$
$=2\mathbb{E}\bigg{[}\frac{1-A}{\pi_{1}(X)}-1\bigg{|}X\bigg{]}.$
@@
$\hat{\pi}_{1}$ $=\arg\min_{\pi_{1}}\max_{h_{1}}\mathbb{E}\left[\left\{\frac{1-A}{\pi_{1}(X)}-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]$
$=\arg\min_{\pi_{1}}\mathbb{E}\left[2\left\{\frac{1-A}{\pi_{1}(X)}-1\right\}\mathbb{E}\bigg{[}\frac{1-A}{\pi_{1}(X)}-1\bigg{|}X\bigg{]}-\mathbb{E}\bigg{[}\frac{1-A}{\pi_{1}(X)}-1\bigg{|}X\bigg{]}\right]$
$=\arg\min_{\pi_{1}}\mathbb{E}\left[\mathbb{E}\bigg{[}\frac{1-A}{\pi_{1}(X)}-1\bigg{|}X\bigg{]}^{2}\right],$
which since zero is achievable by $p(A=0\mid X)$, it implies that
$\mathbb{E}\bigg{[}\frac{1-A}{\hat{\pi}_{1}(X)}-1\bigg{|}X\bigg{]}=0.$
@@
$\mathbb{E}\left[\left\{\frac{1-A}{\hat{\pi}_{1}(X)}-1\right\}h_{1}(X)\right]=\mathbb{E}\left[\mathbb{E}\bigg{[}\frac{1-A}{\hat{\pi}_{1}(X)}-1\bigg{|}X\bigg{]}h_{1}(X)\right]=0.$
@@
    For all $h_{1}\in\mathcal{H}_{1}$, we have
    $\mathbb{E}\left[\left\{\frac{1-A}{\pi_{1}(X)}-1\right\}h_{1}(X)\right]=\mathbb{E}\left[\left\{\frac{\mathbb{E}[1-A\mid X]}{\pi_{1}(X)}-1\right\}h_{1}(X)\right]=0.$
@@
    For all $h_{2}\in\mathcal{H}_{2}$, we have
    $\mathbb{E}\left[\left\{\frac{A\pi_{2}(M,X)}{1-\pi_{1}(X)}-\frac{1-A}{\pi_{1}(X)}\right\}h_{2}(M,X)\right]$
    $=\mathbb{E}\left[\left\{\frac{p(A=1\mid M,X)\frac{p(M\mid A=0,X)}{p(M\mid A=1,X)}}{p(A=1\mid X)}-\frac{p(A=0\mid M,X)}{p(A=0\mid X)}\right\}h_{2}(M,X)\right]$

### v2

@@ 5 Theoretical Analysis
We start by making the observation that the true nuisance functions $\pi_{1}$ and $\pi_{2}$ satisfy conditions (1) and (2) with equality.
@@
    For the true nuisance function $\pi_{1}$, we have
    $\mathbb{E}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)\right]=0,~{}~{}~{}~{}~{}~{}\forall h_{1}\in L^{2}(P_{0,X}).$
@@
    For the true nuisance functions $\pi_{1}$ and $\pi_{2}$, we have
    $\mathbb{E}\left[\left\{A\pi_{2}(M,X)-(1-A)\pi_{1}(X)\right\}h_{2}(M,X)\right]=0,~{}~{}~{}~{}~{}~{}\forall h_{2}\in L^{2}(P_{0,(M,X)}).$
Suppose conditions (1) and (2) hold with equality and $\mathcal{H}_{1}=L^{2}(P_{0,X})$ and $\mathcal{H}_{2}=L^{2}(P_{0,(M,X)})$. Then, it is easy to see that $\hat{\pi}_{1}=\pi_{1}$ and $\hat{\pi}_{2}=\pi_{2}$ with probability one. However, when $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are not the corresponding $L^{2}$ spaces, we can have conditions (1) and (2) hold with equality even though $\hat{\pi}_{1}\neq\pi_{1}$ and $\hat{\pi}_{2}\neq\pi_{2}$. Nevertheless, satisfying conditions (1) and (2) can be sufficient for controlling the bias of the final estimator of the parameter of interest. We formalize this argument by looking at the structure of the bias of $\hat{\psi}^{mr}$.
@@
The bias of the estimator $\hat{\psi}^{mr}$ satisfies the following equality:
$\mathbb{E}[\hat{\psi}^{mr}]-\psi_{0}$ $=\mathbb{E}\left[\left\{(1-A)\hat{\pi}_{1}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\left[\left\{A\hat{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right].$
From the structure of the bias of the estimator $\hat{\psi}^{mr}$, it is evident that if there exists $h_{1}\in\mathcal{H}_{1}$ which well-approximates $\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}$, and there exists $h_{2}\in\mathcal{H}_{2}$ which well-approximates $\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}$, and further conditions (1) and (2) hold, then we can eliminate the bias of the estimator $\hat{\psi}^{mr}$ even if we do not have a perfect fit of $\pi_{1}$ and $\pi_{2}$. This is in fact the main motivation of our two-stage approach. It substitutes fitting $p(A=1\mid X)$ and $p(M\mid A,X)$, which can be a difficult task, by a much weaker requirement based on balancing.
Importantly, under the following positivity assumption, our approach retains the multiple-robustness property, as stated in Proposition 6.
Assumption 2.
For some $\epsilon>0$, we have $p(A=0\mid X)>\epsilon$ and $p(A=1\mid M,X)>\epsilon$.
@@
Under Assumption 2 , $\hat{\psi}^{mr}$ is asymptotically unbiased if at least one of the following conditions holds:.
  • $\|\hat{\pi}_{1}(X)-\pi_{1}(X)\|_{2}=o_{p}(1),$ and $\|\hat{\pi}_{2}(M,X)-\pi_{2}(M,X)\|_{2}=o_{p}(1)$;
  • $\|\hat{\mu}_{1}(M,X)-\mu_{1}(M,X)\|_{2}=o_{p}(1),$ and $\|\hat{\mu}_{2}(X)-\mu_{2}(X)\|_{2}=o_{p}(1)$;
  • $\|\hat{\pi}_{1}(X)-\pi_{1}(X)\|_{2}=o_{p}(1),$ and $\|\hat{\mu}_{1}(M,X)-\mu_{1}(M,X)\|_{2}=o_{p}(1)$.
Figure 2 summarizes the result in Proposition 6. Note that our estimator $\hat{\psi}^{mr}$ is based on an influence function, which always leads to a second-order bias.
@@ 5.1 Asymptotic Normality
    $\|(\pi_{1}(X)-\hat{\pi}_{1,\ell}(X))^{2}\|_{2}=o_{p}(1)$ and $\|(\pi_{2}(M,X)-\hat{\pi}_{2,\ell}(M,X))^{2}\|_{2}=o_{p}(1)$.
  3. (c)
    There exists some constant $C$ such that $\|\mu_{1}^{2}(M,X)\|_{2}$, $\|\mu_{2}^{2}(X)\|_{2}$, and $\mathbb{E}[Y^{4}]$, are upper bounded by $C$.
  4. (d)
    $\big{\|}p(A=0\mid X)\hat{\pi}_{1,\ell}(X)-1\big{\|}_{2}\|\mu_{2}(X)-\hat{\mu}_{2,\ell}(X)\|_{2}=o_{p}(n^{-1/2})$, and
    $\big{\|}p(A=1\mid M,X)\hat{\pi}_{2,\ell}(M,X)-p(A=0\mid M,X)\hat{\pi}_{1,\ell}(X)\big{\|}_{2}\|\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\|_{2}=o_{p}(n^{-1/2})$.
For any choices of the nuisance functions $\{\tilde{\pi}_{1},\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2}\}$, define,
$\Phi(O;\tilde{\pi}_{1},\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2},\psi):=\phi(O;\tilde{\pi}_{1},\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2})-\psi.$
We have the following result regarding the asymptotic behavior of $\hat{\psi}^{mr}$.
@@
Under Assumptions 2 and  3, the estimator $\hat{\psi}^{mr}$ is $\sqrt{n}$-consistent and asymptotically normal and satisfies
$\sqrt{n}(\hat{\psi}^{mr}-\psi_{0})=\sqrt{n}\mathbb{E}_{n}\left[\Phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2},\psi_{0})\right]+o_{p}(1).$
As a corollary of Theorem 2, one can use the influence function of $\psi_{0}$ to obtain confidence intervals for the parameter of interest. For $\ell\in\{1,\dots,L\}$, we estimate the variance of the influence function on fold $I_{\ell}$ as,
$\hat{\sigma}_{\ell}^{2}=\frac{1}{m}\sum_{i=1}^{m}\Phi^{2}(O_{i,\ell};\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell},\hat{\psi}_{\ell}^{mr}),$
where $O_{i,\ell}\in I_{\ell}$, and define,
$\hat{\sigma}^{2}=\frac{1}{L}\sum_{\ell=1}^{L}\hat{\sigma}_{\ell}^{2}.$
Then the $(1-\alpha)$ confidence interval of $\psi_{0}$ can be obtained as,
$\hat{\psi}^{mr}\pm z_{1-\alpha/2}\frac{\hat{\sigma}}{\sqrt{n}},$
where $z_{1-\alpha/2}$ is the $(1-\alpha/2)$-quantile of the standard normal distribution.
@@ 6 Conclusion
In estimating a functional of interest that involves nuisance functions, if one naively aims to perfectly fit the nuisance functions, the resulting estimator of the parameter of interest can be sensitive to even small errors in the nuisance function estimators. Therefore, in order to estimate a functional, it is important to investigate the role of nuisance functions involved in that functional in the structure of the bias. Based on this vision, we proposed a new approach for estimating the nuisance functions of the mediation functional when the influence function-based estimator is used. Our proposed estimators are based on the approach of balancing that can target the bias of the parameter of interest more directly and also can lead to an estimator which is less sensitive to error. We demonstrated that this approach retains the multiple-robustness property. Moreover, we provided conditions on our nuisance function estimators sufficient for consistency and asymptotic normality of the influence function-based estimator of the mediation functional.
@@ Appendix A Proofs
$\arg\max_{h_{1}}\mathbb{E}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]$
$=\arg\min_{h_{1}}\mathbb{E}\left[\left\{(1-A)\pi_{1}(X)-1\right\}^{2}+\frac{1}{4}h^{2}_{1}(X)-\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)\right]$
$=\arg\min_{h_{1}}\mathbb{E}\left[\left(\frac{1}{2}h_{1}(X)-\left\{(1-A)\pi_{1}(X)-1\right\}\right)^{2}\right]$
$=2\mathbb{E}\bigg{[}(1-A)\pi_{1}(X)-1\bigg{|}X\bigg{]}.$
@@
$\breve{\pi}_{1}$ $=\arg\min_{\pi_{1}}\max_{h_{1}}\mathbb{E}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]$
$=\arg\min_{\pi_{1}}\mathbb{E}\left[2\left\{(1-A)\pi_{1}(X)-1\right\}\mathbb{E}\bigg{[}(1-A)\pi_{1}(X)-1\bigg{|}X\bigg{]}-\mathbb{E}\bigg{[}(1-A)\pi_{1}(X)-1\bigg{|}X\bigg{]}^{2}\right]$
$=\arg\min_{\pi_{1}}\mathbb{E}\left[\mathbb{E}\bigg{[}(1-A)\pi_{1}(X)-1\bigg{|}X\bigg{]}^{2}\right],$
which since zero is achievable by $\frac{1}{p(A=0\mid X)}$, it implies that
$\mathbb{E}\bigg{[}(1-A)\breve{\pi}_{1}(X)-1\bigg{|}X\bigg{]}=0.$
@@
$\mathbb{E}\left[\left\{(1-A)\breve{\pi}_{1}(X)-1\right\}h_{1}(X)\right]=\mathbb{E}\left[\mathbb{E}\bigg{[}(1-A)\breve{\pi}_{1}(X)-1\bigg{|}X\bigg{]}h_{1}(X)\right]=0.$
Proof of Proposition 3.
For any give $\hat{\pi}_{1}$, we note that
$\arg\max_{h_{2}}\mathbb{E}\left[\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right]$
$=\arg\min_{h_{2}}\mathbb{E}\left[\left\{A\pi_{2}(M,X)(1-A)\hat{\pi}_{1}(X)\right\}^{2}+\frac{1}{4}h^{2}_{2}(M,X)-\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)\right]$
$=\arg\min_{h_{2}}\mathbb{E}\left[\left(\frac{1}{2}h_{2}(M,X)-\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\right)^{2}\right]$
$=2\mathbb{E}\bigg{[}A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\bigg{|}M,X\bigg{]}.$
Therefore,
$\breve{\pi}_{2}$ $=\arg\min_{\pi_{2}}\max_{h_{2}}\mathbb{E}\left[\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right]$
$=\arg\min_{\pi_{2}}\mathbb{E}\bigg{[}2\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\mathbb{E}\big{[}A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\big{|}M,X\big{]}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\mathbb{E}\big{[}A\pi_{2}(M,X)-(1-A)\pi_{1}(X)\big{|}M,X\big{]}^{2}\bigg{]}$
$=\arg\min_{\pi_{2}}\mathbb{E}\left[\mathbb{E}\bigg{[}A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\bigg{|}M,X\bigg{]}^{2}\right],$
which since zero is achievable by $\frac{p(M,A=0\mid X)\hat{\pi}_{1}(X)}{p(M,A=1\mid X)}$, it implies that
$\mathbb{E}\bigg{[}A\breve{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\bigg{|}M,X\bigg{]}=0.$
Therefore, for all $h_{2}\in\mathcal{H}_{2}$,
$\mathbb{E}\left[\left\{A\breve{\pi}_{2}(M,X)-(1-A)\pi_{1}(X)\right\}h_{2}(M,X)\right]$
$=\mathbb{E}\left[\mathbb{E}\bigg{[}A\breve{\pi}_{2}(M,X)-(1-A)\pi_{1}(X)\bigg{|}M,X\bigg{]}h_{2}(M,X)\right]=0.$
Proof of Proposition 4.
We first show that, for any function $\pi_{1}$, we have
$\max_{h_{1}\in\mathcal{H}}\mathbb{E}_{n}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]-\lambda_{\mathcal{H}_{1}}\|h_{1}\|_{\mathcal{H}_{1}}^{2}$
=$=\{\xi_{n}(\pi_{1})\}^{\top}K_{\mathcal{H}_{1},n}(\frac{1}{n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}\{\xi_{n}(\pi_{1})\},$
where $\xi_{n}(\pi_{1})=\frac{1}{n}\big{(}(1-A_{i})\pi_{1}(X_{i})-1\big{)}_{i=1}^{n}$.
To see this, by representer theorem, the solution to the maximization will be of the form,
$h_{1}(x)=\sum_{j=1}^{n}\alpha^{\ast}_{j}K_{\mathcal{H}_{1}}(X_{j},x),$
and it suffices to find the coefficients $\alpha^{\ast}=(\alpha_{j}^{\ast})_{j=1}^{n}$.
Note that we have,
$\mathbb{E}_{n}[\{(1-A)\pi_{1}(X)-1\}h_{1}(X)]$ $=\sum_{i=1}^{n}h(X_{i})\{\xi_{n}(\pi_{1})\}_{i}={\alpha^{\ast}}^{\top}K_{\mathcal{H}_{1},n}\{\xi_{n}(\pi_{1})\},$
$\mathbb{E}_{n}[h_{1}^{2}(X)]$ $=\frac{1}{n}{\alpha^{\ast}}^{\top}K_{\mathcal{H}_{1},n}^{2}{\alpha^{\ast}},$
$\|h_{1}\|_{\mathcal{H}_{1}}^{2}$ $={\alpha^{\ast}}^{\top}K_{\mathcal{H}_{1},n}{\alpha^{\ast}}.$
Therefore we have,
$\mathbb{E}_{n}\Big{[}\{(1-A)\pi_{1}(X)-1\}h_{1}(X)-\frac{1}{4}h_{1}^{2}(X)\Big{]}-\frac{1}{4}\lambda_{\mathcal{H}_{1}}\|h_{1}\|_{\mathcal{H}_{1}}^{2}$
$={\alpha^{\ast}}^{\top}K_{\mathcal{H}_{1},n}\{\xi_{n}(\pi_{1})\}-{\alpha^{\ast}}^{\top}(\frac{1}{4n}K_{\mathcal{H}_{1},n}^{2}+\frac{1}{4}\lambda_{\mathcal{H}_{1}}K_{\mathcal{H}_{1},n}){\alpha^{\ast}}.$
Taking the derivative and setting it to zero, the optimal coefficients are obtained as,
$\alpha^{\ast}=2(\frac{1}{n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}\{\xi_{n}(\pi_{1})\},$
and we have,
$\max_{h_{1}\in\mathcal{H}_{1}}\mathbb{E}_{n}\Big{[}\{(1-A)\pi_{1}(X)-1\}h_{1}(X)-\frac{1}{4}h_{1}^{2}(X)\Big{]}-\frac{1}{4}\lambda_{\mathcal{H}_{1}}\|h_{1}\|_{\mathcal{H}_{1}}^{2}$
=$=\{\xi_{n}(\pi_{1})\}^{\top}K_{\mathcal{H}_{1},n}(\frac{1}{n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}\{\xi_{n}(\pi_{1})\}.$
Therefore, the outer minimization problem is reduced to:
$\hat{\pi}_{1}=\arg\min_{\pi_{1}\in\Pi_{1}}\{\xi_{n}(\pi_{1})\}^{\top}\Gamma_{1}\{\xi_{n}(\pi_{1})\}+\lambda_{\Pi_{1}}\|\pi_{1}\|_{\Pi_{1}}^{2},$
where $\Gamma_{1}=K_{\mathcal{H}_{1},n}(\frac{1}{n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}$
By representer theorem, the solution to the maximization will be of the form,
$\pi_{1}(x)=\sum_{j=1}^{n}\alpha_{j}K_{\Pi_{1}}(X_{j},x),$
and it suffices to find the coefficients $\alpha=(\alpha_{j})_{j=1}^{n}$.
Note that,
$\xi_{n}(\pi_{1})$ $=\frac{1}{n}(diag(1-A)K_{\Pi_{1},n}\alpha-e_{n}),$
$\|\pi_{1}\|_{\Pi_{1}}^{2}$ $=\alpha^{\top}K_{\Pi_{1},n}\alpha.$
Therefore, we have,
$\min_{\pi_{1}\in\Pi_{1}}\{\xi_{n}(\pi_{1})\}^{\top}\Gamma_{1}\{\xi_{n}(\pi_{1})\}+\lambda_{\Pi_{1}}\|\pi_{1}\|_{\Pi_{1}}^{2}$
$=\min_{\alpha\in\mathbb{R}^{n}}\frac{1}{n^{2}}\alpha^{\top}K_{\Pi_{1},n}diag(1-A)\Gamma_{1}diag(1-A)K_{\Pi_{1},n}\alpha+\frac{2}{n^{2}}e_{n}^{\top}\Gamma_{1}diag(1-A)K_{\Pi_{1},n}\alpha+\lambda_{\Pi_{1}}\alpha^{\top}K_{\Pi_{1},n}\alpha+c,$
which is solved by
$\alpha=-\Big{(}K_{\Pi_{1},n}diag(1-A)\Gamma_{1}diag(1-A)K_{\Pi_{1},n}+n^{2}\lambda_{\Pi_{1}}K_{\Pi_{1},n}\Big{)}^{\dagger}K_{\Pi_{1},n}diag(1-A)\Gamma_{1}e_{n}.$
The derivation for $\beta$ is similar and omitted for brevity.
@@
    For all $h_{1}\in L^{2}(P_{0,X})$, we have
    $\mathbb{E}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)\right]=\mathbb{E}\left[\left\{\mathbb{E}[1-A\mid X]\pi_{1}(X)-1\right\}h_{1}(X)\right]=0.$
@@
    For all $h_{2}\in L^{2}(P_{0,(M,X)})$, we have
    $\mathbb{E}\left[\left\{A\pi_{2}(M,X)-(1-A)\pi_{1}(X)\right\}h_{2}(M,X)\right]$
    $=\mathbb{E}\left[\left\{\frac{p(A=1\mid M,X)p(M\mid A=0,X)}{p(M,A=1\mid X)}-\frac{p(A=0\mid M,X)}{p(A=0\mid X)}\right\}h_{2}(M,X)\right]$